Accounts Receivables Valuation (under allowance method)

The ledger of Meteo Company at the end of the current year shows Accounts Receivable $420,000, Sales Revenue $920,000, and Sales Returns and Allowances $10,000.

Instructions: If Allowance for Doubtful Accounts has a credit balance of $3,000 in the trial balance, journalize the adjusting entry at December 31, assuming bad debts are expected to be

(1) 1.5% of net sales, and

(2) 8% of accounts receivable.

(c) If Allowance for Doubtful Accounts has a debit balance of $600 in the trial balance, journalize the adjusting entry at December 31, assuming bad debts are expected to be (1) 0.85% of net sales and (2) 6.5% of accounts receivable.
Answer #1

net sales = sales - sales returns

net sales = 920,000 - 10,000 = 910,000

1.5% of net sales = 13650

8% of accounts receivable = 420,000*8% =33600

total of doubtful debt = $47250

available balance is $3000

balance is $44250

bad debt expense account debit $44250

to allowance for doubtful account $44250

2.

0.85% net sales = 910,000 * 0.85% = $7735

6.5% of AR = 420,000 * 6.5% =27300

total = $35035

available balance is 600 debit

balance = 35035 + 600 = $35635

journal entry

bad debt expense account debit $35635

to allowance for doubtful account $35635